### v5

# Practical solutions to the relative pose of three calibrated cameras

Abstract
We study the challenging problem of estimating the relative pose of three calibrated cameras from four point correspondences. We propose novel efficient solutions to this problem that are based on the simple idea of using four correspondences to estimate an approximate geometry of the first two views. We model this geometry either as an affine or a fully perspective geometry estimated using one additional approximate correspondence. We generate such an approximate correspondence using a very simple and efficient strategy, where the new point is the mean point of three corresponding input points. The new solvers are efficient and easy to implement, since they are based on existing efficient minimal solvers, i.e., the 4-point affine fundamental matrix, the well-known 5-point relative pose solver, and the P3P solver. Extensive experiments on real data show that the proposed solvers, when properly coupled with local optimization, achieve state-of-the-art results, with the novel solver based on approximate mean-point correspondences being more robust and accurate than the affine-based solver.
1 Introduction
00footnotetext: ∗ Equal contributionCamera geometry estimation is crucial in many computer vision applications, e.g., visual navigation [DBLP:journals/ram/ScaramuzzaF11], Structure-from-Motion [Snavely-IJCV-2008], augmented reality [Castle08ISWC], self-driving cars [hane20173d], and visual localization [Sattler16PAMI]. Due to noise and outliers in input correspondences, the predominant way for camera geometry estimation is to use a hypothesis-and-test framework, e.g., RANSAC [Fischler-Bolles-ACM-1981, Chum-2003, DBLP:journals/pami/RaguramCPMF13, barath2017graph]. For RANSAC-like methods, using as few (ideally the minimal number of) correspondences as possible for estimation is important since the number of RANSAC iterations (and thus its run-time) grows exponentially with the number of correspondences required for the model estimation.
Minimal camera geometry problems often result in complex systems of polynomial equations. Efficient algebraic methods helped solve many previously unsolved problems [Stewenius-CVPR-2005, bujnak_cvpr2008, larsson2019revisiting, DBLP:conf/cvpr/KukelovaP07, kukelova2013real, Stewenius-ISPRS-2006]. Still, they fail to generate efficient and/or numerically stable solutions for some configurations. In this paper, we study one such challenging problem: Estimating the relative pose of three calibrated cameras. This problem has received attention for a long time [holt1995, Quan2006, leonardos_cvpr2015, Martyushev16, Aholt2014]. However, due to its complexity, it is still not considered fully solved. There are no efficient and practical solutions for most of the minimal configurations of point and/or line correspondences [Kileel2017]. One such configuration that is particularly interesting is the notoriously difficult configuration of four points in three calibrated views [Quan2006, Nister-5pt-PAMI-2004], known as the 4p3v problem.
State-of-the-art algebraic and numerical methods are known to fail in generating efficient and numerically stable solutions to the 4p3v problem. The existing methods for solving this problem are only approximate [Hruby_cvpr2022, DBLP:journals/ijcv/NisterS06]. By solving only for one (or a few) solutions from the 272 solutions of the 4p3v problem [Hruby_cvpr2022], and by discretely sampling the space of potential solutions [DBLP:journals/ijcv/NisterS06], the existing 4p3v methods can often fail, i.e., the returned solution can be, in general, arbitrarily far from the geometrically correct solution. To decrease the failure rate, both methods [Hruby_cvpr2022, DBLP:journals/ijcv/NisterS06] require a lot of tuning and are challenging to re-implement.111There is no publicly available implementation for [DBLP:journals/ijcv/NisterS06]. Implementation of [Hruby_cvpr2022] is quite complex and requires a non-negligible effort to run.
In contrast to complex solutions to the configuration of four points in three views, there is a simple and efficient solution for the configuration where five correspondences are detected in the first two views and three of these points are also visible in the third view. In this case, the 5pt+P3P solver first estimates the relative pose of two cameras from five correspondences using an efficient 5pt solver [Nister-5pt-PAMI-2004], and then registers the third camera using a P3P solver [lambda-twist].
Following the idea of first estimating the relative pose of two cameras, we propose two novel approaches for solving the 4p3v problem. Our solutions are based on the simple idea of using four correspondences to estimate an approximate geometry of the first two views and then registering the third camera using a P3P solver [lambda-twist].
In the first approach, we approximate the geometry of the first two views using affine cameras. The approximation using affine cameras was shown to provide good accuracy for relative pose estimation if coupled with geometry refitting and local optimization inside RANSAC [pritts_ivcnz13]. We denote the proposed affine-based 4p3v solver as 4p3v(A).
In the second approach, we use a less restrictive approximation. In this case, we estimate the full perspective geometry, i.e., the full 5DoF essential matrix, using four input correspondences and one additional approximate correspondence in two views. We generate the approximate correspondence using the locations of three of the four input correspondences. Under the assumption of a para-perspective projection, i.e., of affine geometry, the mean point of three 3D points is projected to the mean points in both images [Zhang2014]. Thus, the new correspondence is generated as the correspondence between the mean points of three corresponding points detected in two views. This approximate mean-point correspondence can also be seen as a correspondence under the -order approximation of the homography defined by the plane passing through the three corresponding 3D points. We denote the proposed mean point-based 4p3v solver as 4p3v(M).
While the proposed 4p3v(M) solver returns significantly more accurate poses than the 4p3v(A) solver, both solvers may not be accurate enough if not properly treated inside RANSAC. Thus, in this paper, we propose several ways to improve the accuracy of the proposed approximate solvers. (i) To compensate for noise in the mean point correspondences, we introduce the 4p3v(M) solver that generates two additional correspondences in the vicinity of the mean point, and in the first step calls the 5pt solver [Nister-5pt-PAMI-2004] three times. (ii) To improve the approximate geometry estimated between the first two views, we refit this geometry using a non-minimal relative pose solver and inliers obtained from the approximate geometry (ENM). (iii) We use the fourth correspondence in the third view to filter out geometrically infeasible solutions (+F) and to refine the solutions on four input correspondences in three views using just a few iterations of Levenberg-Marquardt (LM) refinement (+R). While conceptually simple and efficient, the novel solvers achieve state-of-the-art results on real data.
The contributions of the paper are as follows:
We propose two groups of novel solutions for the well-known and challenging 4p3v problem: 4p3v(A)- and 4p3v(M)-based solvers. These solutions are based on the simple idea of using four correspondences to estimate an approximate geometry in the first two views. Compared to state-of-the-art 4p3v solvers [Hruby_cvpr2022, DBLP:journals/ijcv/NisterS06], which are non-trivial and difficult to re-implement such that they are numerically stable and fast, our new solutions can be easily implemented using existing efficient implementations of the linear 4pt affine fundamental matrix solver [hartley2006multiple], the 5pt solver [Nister-5pt-PAMI-2004] and the P3P solver [lambda-twist]. The source code is available at https://github.com/kocurvik/threeview and https://doi.org/10.5281/zenodo.16599943. The data used are available and can be accessed at https://doi.org/10.5281/zenodo.16603086.
We present several ways of improving the accuracy and the speed of the proposed solvers, as well as a strategy for efficiently using these approximate solvers inside a RANSAC-style paradigm. We show that the new solvers achieve state-of-the-art results in terms of accuracy on real data. While both 4p3v(A)- and 4p3v(M)-based solvers achieve comparable results when coupled with the suggested non-minimal geometry refitting and local optimization, our 4p3v(M)-based solvers are more robust to the scene geometry and RANSAC inlier thresholds.
To our knowledge, we are the first to extensively evaluate solutions to the 4p3v problem on a large variety of real-world scenes and within state-of-the-art RANSAC frameworks, and to compare them to the baseline 5pt+P3P solver. We report results on 3 datasets, consisting of 18 different scenes and altogether 90,000 camera triplets.
2 Related work
Estimating the relative pose of three cameras from a minimal number of point and line correspondences is known as an extremely challenging problem. For three uncalibrated cameras, 6 point correspondences are necessary to estimate the trifocal tensor, with a solution known for a long time [Quan_pami95, Torr97a]. Solutions to three minimal combinations of points and lines are presented in [Oskarsson_bmvc2004]. The configuration of 9 lines is more challenging and was solved only recently [larsson2017efficient]. Yet, the solver is far from practical and runs 17.8s.
For calibrated cameras, the configuration that attracts most of the attention is the configuration of four points in three views (the 4p3v problem). Note that this is not a minimal configuration since it generates 12 constraints for 11 degrees-of-freedom (DoF). The 4p3v problem is known to be extremely difficult to solve. Several papers present mostly theoretical results [leonardos_cvpr2015, Martyushev16, Aholt2014]. For four triplets of exact points without noise, it is shown that the 4p3v problem has, in general, a unique solution [holt1995, Quan2006].
To the best of our knowledge, there are only two reasonably efficient solutions to the 4p3v problem reported in the literature. The first solver [DBLP:journals/ijcv/NisterS06] is based on a similar idea as our solvers, i.e., to first estimate the relative pose of two cameras and then register the third camera using a P3P solver [lambda-twist]. To compute the pose of the first two cameras using information only from four point correspondences, the solver needs additional information about the position of one epipole. The paper shows that the four point correspondences between two calibrated views constrain the epipole in each image to lie on a curve of degree ten. Thus, the solver performs one-dimensional exhaustive search and sweeps a -degree curve of possible epipoles. For each potential epipole, it computes the relative pose of two cameras, registers the third camera using three triangulated points, and finally extracts the solution minimizing the reprojection error of the fourth point in the third view. Evaluation of the solver on one potential epipole is fast. Yet, in contrast to our proposed 4p3v(M)-based solvers, the error of the sampled epipole for one fixed point on the curve is not bounded since the true epipole can lie anywhere on the curve, and in many scenarios it is very far from the image center (e.g., outside the image). Thus, to obtain reasonably accurate and stable results, usually 1,000 candidates need to be evaluated. Even then, refinement at multiple local minima is required to improve the accuracy. The runtimes reported for this solver were depending on the number of points searched. On a small number of synthetic experiments, the paper shows that the translation error returned by the proposed solver is usually higher than the error returned by the 5pt+P3P solver. The literature does not compare against [DBLP:journals/ijcv/NisterS06] as there is no publicly available implementation and it is hard to re-implement.
The second efficient solver to the 4p3v problem was published only recently [Hruby_cvpr2022]. In this paper, the authors first transform the 4p3v problem into a minimal problem by considering a line passing through the last correspondence in the third view. The resulting system of equations is solved using an efficient Homotopy continuation (HC) method [Fabbri_CVPR2020, SommeseAndrewJ2005Tnso]. To avoid computing large numbers of spurious solutions, an MLP-based classifier is trained. For a given problem , it selects one or several starting problem-solution pairs (so-called anchors), such that the geometrically meaningful/correct solution of can be obtained by HC starting from this anchor. This strategy is fast, running on average per solution. However, it has a high failure rate. The success rate of the 4p3v solver reported in [Hruby_cvpr2022] on two test datasets and data without noise is . [Hruby_cvpr2022] do not show results for a real scenario, i.e., a RANSAC-like framework with noisy data. Providing such an evaluation, we show that our much simpler solvers outperform [Hruby_cvpr2022].
Solutions to the 4p3v problem for orthographic and para-perspective views were presented in [xu_ortho4p3v, Higgins-4p3v91]. In [Higgins-4p3v91], the author suggested an iterative approach for updating to perspective views, but reported results only on a few synthetic instances. According to our extensive experiments, the update does not work on real data with general perspective cameras. This solver is returning large errors even after incorporating it into RANSAC with local optimization. We see two main reasons: (i) The set of inliers that satisfy an approximate para-perspective camera model in all three images is usually quite small, and (ii) the estimated model is often very far from the perspective optimum, and thus local optimization does not converge to a good solution.
In [Duff_PL1P, Kileel2017, duff2019plmp], the authors aim to classify and derive the number of solutions for different minimal configurations of points and lines in three calibrated views. Solutions to two minimal configurations combining points and lines were proposed in [Fabbri_CVPR2020] and solved using a HC method [SommeseAndrewJ2005Tnso]. Due to their complexity, the solvers are not practical. A GPU HC method was also used to solve minimal problems of four points/six lines in three views for a generalized camera in [Ding_2023_ICCV] and of four points in three cameras with an unknown shared focal length (4p3vf) in [Chien_2022_CVPR, Cin_2024_CVPR]. Efficient GPU implementations of the 4p3vf solvers run to . These times are still too slow for practical applications.
An approximation of perspective cameras using affine ones was used in several papers on camera geometry estimation [Collins2014InfinitesimalPP]. [oberkampf1996] and [Horaud97] propose iterative approaches to the absolute camera pose estimation problem, i.e., the PnP problem. Both methods first compute the pose for an affine camera (weak- or para-perspective). Then the error induced by the affine camera approximation is used to adjust the constraints on the pose and recompute it. These methods do not guarantee convergence to the perspective model solution. In [pritts_ivcnz13], affine cameras are used to efficiently solve the relative pose problem of two uncalibrated cameras from two affine correspondences. These affine correspondences are transformed to six point or two ellipse correspondences and then used to estimate the affine fundamental matrix. The paper also proposes a RANSAC framework that uses local optimization [lebedaLO] to estimate the full fundamental matrix using inliers from the approximate affine model.
Points that are sampled based on feature geometry to generate point correspondences from affine or scale- and orientation-invariant feature correspondences were also used in [perd2006epipolar, BEEP, pritts-cvpr2018]. In contrast, our M-based solvers use only point correspondences, without associated feature geometry, to generate an additional point correspondence.
3 Estimating the relative pose of three cameras
In this section, we describe different solutions for estimating the relative pose of three calibrated cameras. We start with a baseline solution for the minimal configuration of three points visible in all three cameras and two additional points visible in two of the three cameras (the (5,5,3) configuration). Next, we present our novel solutions for the configuration of four points visible in all three cameras (the (4,4,4) configuration). This configuration generates an over-constrained problem. In this case, we have one more constraint than DoF. A minimal solution would need to drop one constraint, e.g., by considering only a line passing through one of the points in the third view [Hruby_cvpr2022] or by considering a “half” point correspondence. Since, in practice, we always have full correspondences and sampling one less point in one view leads to an under-constrained problem, the (4,4,4) configuration, is usually considered “minimal”.
5pt+P3P solver: The 5pt+P3P solver first estimates the relative pose of two cameras from 5 image point correspondences using the efficient 5pt solver [Nister-5pt-PAMI-2004]. Next, the three points visible in all three views are triangulated. Finally, the third camera is registered using the three 2D-3D point correspondences and the well-known efficient P3P solver [lambda-twist].
This straightforward solver, which is based on existing efficient solvers [Nister-5pt-PAMI-2004, lambda-twist], was discussed in several works [Duff_PL1P, Nister-5pt-PAMI-2004, Rodehorst2017, DBLP:journals/ijcv/NisterS06]. [DBLP:journals/ijcv/NisterS06] showed that the 5pt+P3P solver performs better than their dedicated 4p3v solver on synthetic data. Yet, the most recent works [Fabbri_CVPR2020, Hruby_cvpr2022] that study the three view relative pose problem do not discuss the 5pt+P3P solver and do not use it as a baseline for comparison. To the best of our knowledge, the performance of this solver on real data and within state-of-the-art RANSAC frameworks in the context of the 4p3v problem has not been extensively studied. Our paper fills this gap in the literature.
3.1 Approximate solutions to the 4p3v problem
In contrast to the (5,5,3) configuration, the (4,4,4) configuration, i.e., the configuration of four points in three views, leads to significantly more complex equations. State-of-the-art algebraic and numerical methods are known to fail in generating efficient and numerically stable solutions to these equations. Thus, solutions to the 4p3v problem require some approximations to be practical in real-world applications. This section introduces several practical approximate solutions to the 4p3v problem. Similarly to the 5pt+P3P solver, we decompose the problem into the problem of first estimating the relative pose of two cameras and then registering the third camera with an efficient P3P solver [lambda-twist]. The idea is to assume an approximate geometry only in the first two views and use four input point correspondences to efficiently estimate this geometry. Next we describe two groups of such solvers.
4p3v(A) solver: The first group of solvers approximates the geometry in the first two views using affine cameras. Although the approximation can be quite far from the correct geometry, [pritts_ivcnz13] showed that an approximate affine camera model provides good accuracy for relative pose estimation if coupled with geometry refitting (using the full 7DoF fundamental matrix) and local optimization inside RANSAC. The proposed 4p3v(A) solver first uses four point correspondences in two views to efficiently estimate the affine fundamental matrix [hartley2006multiple] and then registers the third camera using three triangulated points and a P3P solver [lambda-twist].
4p3v(M) solver: The affine camera model used in the 4p3v(A)-based solvers can be quite imprecise. This motivates us to introduce solvers that are based on a less restrictive approximation. In this case, we approximate only one point correspondence in two views. Under the assumption of a para-perspective projection, i.e., of affine geometry, it is known that the mean point of three 3D points is projected to the mean points of their projections in both images [Zhang2014]. Thus, we generate one new approximate correspondence in two views as the correspondence between the mean points of three corresponding points detected in these views. Let be the mean point of three points , in the view , then is our new approximate correspondence. The new 4p3v(M) solver solves the 4p3v problem by first estimating the full 5DoF essential matrix from four original correspondences and the correspondence . This is done using the efficient 5pt relative pose solver [Nister-5pt-PAMI-2004]. As with previous solvers, the third camera is registered using a P3P solver [lambda-twist]. In short, the 4p3v(M) solver solves the 4p3v problem using the 5pt+P3P solver and one approximate correspondence.
As we show on large amounts of synthetic and real data (see Sec. 4 and Supp. mat. (SM)), the 4p3v(M) solver provides much more accurate estimates than the 4p3v(A) solver. Moreover, even in the basic form (without the modifications presented in the next section), on many scenes it performs comparably to the state-of-the-art HC solver [Hruby_cvpr2022]. This can be attributed to several facts and observations: (1) The correspondence does not need to be seen as a correspondence of points that are projections of the mean point of three 3D points and (in which case both and would have some error). We can look at this correspondence as a correspondence in which we fix a point in one view, e.g., , and generate a corresponding point in the second view. In this case, to generate a good correspondence, we only require the point in the second view to be reasonably close to the epipolar line defined by the mean point in the first view, i.e., the 2D point does not need to correspond to one particular 3D point with a given depth. (2) The ray from the center of the first camera through the mean point intersects the plane of and at point , which lies inside the triangle formed by these 3D points. The projection of into the second camera lies inside the triangle formed by . By construction, this projection lies on the epipolar line . Thus the epipolar line defined by passes through the triangle defined by . Consequently, the maximum distance of in the second image from the epipolar line is bounded by the maximum distance of from . For a formal lemma and proof, see SM. (3) For practical applications, when used in RANSAC, it is not necessary that each triplet generates a good correspondence . Samples with a high level of noise in the mean-point correspondence are filtered inside RANSAC.222This property was also used in the HC solver [Hruby_cvpr2022], which completely fails for many samples. These samples are filtered within RANSAC. On a large number of different scenes, we observed that even if some image pairs have triplets of points that generate very noisy mean-point correspondences, there are usually enough triplets for which the noise in is reasonably small to lead to good estimates. (4) Four point correspondences in two views usually fix the space of possible poses such that the correspondence, even if noisy, often generates a pose that is not very far from the ground truth pose. Such a pose is usually sufficient for filtering out outliers and a good initialization for non-linear optimization on the original four points in three views and subsequent local optimization on detected inliers. We support our observations by experiments on a large amount of data.
3.2 Making approximate solvers practical
The 4p3v(A)-based solvers, when used without any modifications, generally provide imprecise results even when used inside RANSAC with local optimization on three views. While the accuracy of the pure 4p3v(M) solver inside LO-RANSAC [lebedaLO] (RANSAC with Local Optimization) is much better, there is still room for improvement. Here we present several simple modifications of these solvers that significantly boost their performance.
4p3v(M) solver: The mean point correspondence used in the 4p3v(M) solver can provide a good approximation of a correct correspondence. Yet, as mentioned above, it can also be noisy. In the 4p3v(M) solver, we thus, in addition to the mean point of three points in the second image, generate two additional points. These points are (1) if the longest dimension of the triangle is in the x-direction or (2) if it is in the y-direction. All three points, i.e., , , and are placed in correspondence with the mean point . The 4p3v(M) solver in the first step calls the 5pt solver [Nister-5pt-PAMI-2004] three times, with the correspondence being either , , or . The results of these three 5pt solvers are collected to create hypotheses for the relative pose of the first two cameras inside RANSAC. The shift is selected relative to the size of the triangle .
Early non-minimal refitting (ENM): The geometry estimated between the first two cameras using the proposed 4p3v(A) and 4p3v(M)- based solvers is only approximate. In standard LO-RANSAC, such a geometry is optimized in local optimization (LO) after registering the third view. However, this can lead to propagation of errors into triangulated points and subsequently into errors in the pose of the third camera. Inliers in three views w.r.t. such cameras together with imprecise pose initializations may not be sufficient for LO to converge to a good solution. We observed this especially for 4p3v(A)-based solvers. Fortunately, the design of our solvers allows us to optimize the geometry already after estimating the approximate geometry between the first two cameras. Here, even a very imprecise geometry returned by the 4p3v(A) solvers is usually sufficient to filter out outliers. Thus, after running the 4pt /5pt solver in the first step of the proposed solvers, we use the estimated approximate models to detect inliers in two views. We then refit the estimated geometry using the non-minimal version of the 5pt solver [Nister-5pt-PAMI-2004], which instead of a 4-dim null space of a matrix uses the last four vectors from the SVD/QR decomposition of a matrix.
point in the third view: 4p3v(A) and 4p3v(M) solvers actually solve the configuration (4,4,3), i.e., they do not use the information from the point in the third view. The information from can be used in two different ways:
Filtering (+F): can be used to filter out geometrically infeasible solutions returned by the proposed solvers. Note that the P3P and 5pt solvers used inside the proposed methods and the -based strategy return multiple solutions that can be evaluated w.r.t. to filter out spurious solutions. Since the returned solutions can be affected by the proposed approximation, we do not simply select the solution with the smallest error on , but we keep all solutions that have an epipolar error on smaller than twice the threshold used inside RANSAC. Our experiments show that this filtering can improve the speed of the proposed solvers. However, as a trade-off, there is sometimes a small drop in the accuracy of the solvers since, in some cases, geometrically correct solutions are filtered out.
Refinement (+R): can be used to refine the solutions returned by the proposed solvers. We want solutions that minimize the epipolar error on the original 4 points in all 3 views, i.e., that solve the original (4,4,4) configuration. (4,4,4) is an overconstrained configuration; thus, for noisy data, there is, in general, no solution with zero error on all 4 points in 3 views. We refine the poses by minimizing the epipolar error of the original four points in three views using LM optimization, initialized using the solutions from the 4p3v(A)- and 4p3v(M)-based solvers. Experiments with different numbers of iterations show that two iterations are usually sufficient to obtain an improvement (see SM).
4 Experiments
We extensively evaluated the proposed solvers on a large variety of synthetic and real data to test their robustness to noise, outliers, and scene properties, and to assess their performance inside state-of-the-art RANSAC-frameworks [barath2017graph, PoseLib]. We compare our novel solvers with the homotopy continuation 4p3v(HC) solver [Hruby_cvpr2022] and the 5pt+P3P baseline minimal solver for the (5,5,3) configuration.
Experimental setup. To obtain feature correspondences, we use SuperPoint [detone2018superpoint] features with the LightGlue [lindenberger2023lightglue] matcher. We extract at most 2048 features per image. We perform matching for all three image pairs and keep only those matches that were consistently matched across all three views. We perform evaluation within two RANSAC frameworks: PoseLib [PoseLib] and GC-RANSAC [barath2017graph]. For the 5pt solver, we use [Nister-5pt-PAMI-2004] and for the P3P solver, we use [lambda-twist]. In PoseLib, we perform LO [Chum-2003] using LM optimization. In GC-RANSAC, we perform LO using non-minimal solvers [Nister-5pt-PAMI-2004, dlspnp] for fitting models to larger-than-minimal samples. We evaluated different shifts for our -based solvers using a validation scene (for the ablation study, see SM) and selected . The choice of is not critical, i.e., different values perform similarly.
Evaluation measures. Inspired by [IMC2020], we define the pose error as , where and are the angular errors of rotation and translation for pair in degrees [IMC2020]. We also report AUC values [IMC2020] at different thresholds for the pose error. We include results for an alternative pose error definition which includes and in SM.
Approximate camera geometry. The first experiments aim to support our idea of estimating approximate geometry in the first two views. The accuracy of the used approximations depends on a large number of variables, including the depths of the points w.r.t. the cameras, the angle under which the points are observed, the type of motion, etc. A detailed analysis of all these factors, e.g., through synthetic experiments, is beyond the scope of this paper.
An approximation error introduced by the para-perspective projection, i.e., the affine geometry, and its effect on absolute camera pose determination is studied in the literature [Zhang2014, Horaud97]. Here, we thus only study the effect of the proposed approximation on relative pose estimation for real-world data and interesting synthetic scenarios. In the following experiments, all solvers are tested on two views without RANSAC. The two-view variants of our solvers are denoted as 4p(A), 4p(M), and 4p(M).
In the first set of experiments, we consider several interesting synthetic setups. Fig. 2 presents the results for increasing the angle between the principal axes of the two cameras. We generate 3D points in a 2000x2000x100 cube uniformly at random. The distances of the camera centers are random between 1000 and 1200. The cameras are looking towards the scene and their principal axes form a specified angle. To simulate realistic scenarios, we add 1px noise to the image correspondences and 20% outliers.333Outliers are only used in the ENM part to simulate that for approximate geometry outliers might be contained in the non-minimal samples. It can be seen that the results for both approximate solvers deteriorate with increasing angle, with 4p(A) being significantly less accurate than 4p(M) and 4p(M). However, ENM can significantly decrease errors, with the 4p(M)+ENM solver returning almost identical results to 5pt+ENM. Results for varying distances of the cameras from the scene, depth of the scene, and image noise are in SM.
Similar observations as on synthetic data can be made on real-world data. Tab. 1 shows results on different scenes from the PhotoTourism dataset [IMC2020]. As can be seen, the approximate solvers are not as accurate as the 5pt solver when used without ENM. The gap between the 4p(M)/4p(M) solvers and the 5pt solver is for some scenes larger (Sacre Coeur, Trevi Fountain), and for some noticeably smaller (Reichstag, St. Peter’s Square), showing that the performance of our solvers is scene-dependent. The 4p(A) solver returns large errors. The accuracy of the 4p(M)/4p(M) solvers is not very far from the 5pt solver, with the 4p(M) solver even slightly outperforming the 5pt solver on some scenes (Reichstag, Taj Mahal). ENM helps to increase the precision of all solvers, especially 4p(A). Still, the 4p(A)+ENM solver returns quite high errors. However, in practice, all solvers are used inside RANSAC, where the current best pose is locally optimized on inliers. LO compensates for less accurate pose estimates and, as we show on real experiments for three as well as two views (see SM), it can suppress the errors of the 4p(A)/4p3v(A)-based solvers, making them comparable to the 4p(M)/4p3v(M)-based solvers. For the three view scenario, the performance of all of our solvers can be further improved by pose refinement (+R) and filtering (+F).
| MED (∘) | perc. (∘) | |||||||
| w/o ENM | ||||||||
| Scene | 5p(E) | 4p(M) | 4p(M) | 4p(A) | 5p(E) | 4p(M) | 4p(M) | 4p(A) |
| Brandenburg Gate | 16.14 | 19.02 | 15.69 | 75.32 | 7.22 | 9.83 | 7.46 | 45.99 |
| Buckingham Palace | 18.87 | 21.12 | 17.74 | 66.19 | 8.04 | 10.01 | 8.23 | 40.59 |
| Colosseum Exterior | 17.40 | 23.50 | 18.93 | 66.67 | 6.44 | 11.36 | 8.34 | 40.82 |
| Grand Place Brussels | 18.40 | 20.28 | 16.78 | 69.31 | 8.04 | 10.09 | 8.07 | 44.45 |
| Notre Dame Front Facade | 15.82 | 23.97 | 19.61 | 71.60 | 6.40 | 12.27 | 9.01 | 48.39 |
| Palace of Westminster | 16.04 | 17.62 | 14.72 | 69.38 | 4.99 | 7.98 | 6.40 | 46.49 |
| Pantheon Exterior | 22.65 | 25.71 | 21.31 | 61.33 | 10.65 | 14.09 | 10.94 | 38.12 |
| Reichstag | 12.15 | 12.99 | 10.27 | 82.85 | 5.15 | 6.16 | 4.57 | 60.12 |
| Sacre Coeur | 11.81 | 17.61 | 14.31 | 74.21 | 3.69 | 7.95 | 5.87 | 43.72 |
| St. Peter’s Square | 17.85 | 18.75 | 15.48 | 72.75 | 8.51 | 9.71 | 7.66 | 44.60 |
| Taj Mahal | 9.95 | 11.20 | 8.56 | 82.74 | 3.77 | 4.90 | 3.63 | 64.27 |
| Temple Nara Japan | 18.57 | 21.53 | 17.01 | 64.97 | 7.60 | 10.63 | 8.16 | 32.21 |
| Trevi Fountain | 20.93 | 27.77 | 22.89 | 50.12 | 8.10 | 13.89 | 10.78 | 31.32 |
| w/ ENM | ||||||||
| Scene | 5p(E) | 4p(M) | 4p(M) | 4p(A) | 5p(E) | 4p(M) | 4p(M) | 4p(A) |
| Brandenburg Gate | 14.53 | 17.05 | 14.38 | 28.40 | 6.74 | 8.86 | 6.93 | 16.96 |
| Buckingham Palace | 16.71 | 18.76 | 16.07 | 35.07 | 7.35 | 8.96 | 7.46 | 19.72 |
| Colosseum Exterior | 16.21 | 21.46 | 17.70 | 34.90 | 6.35 | 10.71 | 8.04 | 21.98 |
| Grand Place Brussels | 16.21 | 17.96 | 15.23 | 31.53 | 7.27 | 8.85 | 7.24 | 18.40 |
| Notre Dame Front Facade | 15.05 | 22.08 | 18.46 | 34.49 | 6.30 | 11.65 | 8.71 | 22.08 |
| Palace of Westminster | 13.83 | 14.95 | 12.74 | 29.06 | 4.68 | 7.13 | 5.84 | 16.41 |
| Pantheon Exterior | 21.19 | 24.03 | 20.30 | 37.21 | 10.32 | 13.28 | 10.55 | 24.42 |
| Reichstag | 9.32 | 10.03 | 8.36 | 17.69 | 4.40 | 5.06 | 3.97 | 9.39 |
| Sacre Coeur | 10.71 | 15.19 | 12.64 | 24.92 | 3.61 | 7.23 | 5.49 | 13.41 |
| St. Peter’s Square | 16.01 | 16.94 | 14.35 | 28.60 | 7.75 | 8.63 | 7.03 | 15.68 |
| Taj Mahal | 8.32 | 9.32 | 7.43 | 15.85 | 3.34 | 4.10 | 3.18 | 7.72 |
| Temple Nara Japan | 16.24 | 19.38 | 15.74 | 29.97 | 6.83 | 9.19 | 7.32 | 16.65 |
| Trevi Fountain | 20.22 | 26.42 | 22.09 | 39.67 | 8.02 | 13.43 | 10.53 | 26.62 |
In the last set of experiments, we study the accuracy of the mean point correspondence. We sample 100 four-tuples of point correspondences consistent with the ground truth relative pose, i.e., inliers, for each image pair in scenes from the PhotoTourism dataset [IMC2020]. We use the first three correspondences to define the triangles in both images. Then, we establish correspondences between the mean of the triangle in one image and various points in the triangle in the second image. We express points in the second triangle via their barycentric coordinates and uniformly sample barycentric coordinates , such that .
Fig. 3 shows the results for the rotation error and the percentage of inliers consistent with the pose obtained with the 4p(M) solver on the St. Peter’s Square scene. The optimum of the metrics is reached around the mean point of the triangles (the mean values of 2D Gaussians fitted to the results are very close to the mean point of the triangles). A detailed description, the translation and the symmetric epipolar errors, and the results for more scenes are in SM. For all tested scenes, we observed a similar behavior.
Experiments on real data. We test the solvers on all scenes from the PhotoTourism dataset [snavely2006photo, IMC2020] which provide ground truth poses and intrinsics via a COLMAP [Schoenberger2016CVPR] reconstruction. In the results, we do not include the St. Peter’s Square scene that we used for the validation of and the number of refinement iterations (see SM). We also include results for the Cambridge Landmarks dataset [kendall2015cambridge] (except the Street scene, which is commonly not used due to issues with its ground truth) and Aachen Day-Night v1.1 [zhang2021aachen]. For PhotoTourism and Aachen, we use the images in their original resolution. For Cambridge Landmarks, we resize them so that the larger side is 800 px. For each scene, we sample 5,000 random image triplets with at least 10 matches obtained with [detone2018superpoint, lindenberger2023lightglue] with at least overlap [IMC2020].
| PhotoTourism [IMC2020] | Cambridge Landmarks [kendall2015cambridge] | Aachen Day-Night v1.1 [zhang2021aachen] | ||||||||||||||||
| Estimator | AVG | MED | AUC@5 | @10 | @20 | Runtime | AVG | MED | AUC@5 | @10 | @20 | Runtime | AVG | MED | AUC@5 | @10 | @20 | Runtime |
| 4p3v(HC) [Hruby_cvpr2022] | 6.37 | 1.62 | 62.41 | 73.53 | 81.94 | 64.10 | 8.73 | 2.67 | 47.99 | 64.58 | 76.65 | 60.11 | 12.43 | 3.55 | 42.28 | 55.20 | 66.55 | 67.26 |
| 5pt+P3P | 5.40 | 1.60 | 62.93 | 74.30 | 83.00 | 33.77 | 7.46 | 2.65 | 48.36 | 65.33 | 77.75 | 24.04 | 10.52 | 3.38 | 43.43 | 56.73 | 68.28 | 53.34 |
| 5pt+P3P+ENM | 5.23 | 1.56 | 63.77 | 75.03 | 83.50 | 48.79 | 7.20 | 2.63 | 48.74 | 65.68 | 78.04 | 34.82 | 10.59 | 3.41 | 43.32 | 56.64 | 68.06 | 71.61 |
| 5pt+P3P+R+F | 5.10 | 1.50 | 64.90 | 75.90 | 84.11 | 38.12 | 7.15 | 2.58 | 49.42 | 66.35 | 78.55 | 31.82 | 10.37 | 3.36 | 43.73 | 57.15 | 68.71 | 54.34 |
| 5pt+P3P+R+F+ENM | 4.99 | 1.50 | 65.18 | 76.20 | 84.30 | 56.52 | 6.98 | 2.58 | 49.38 | 66.31 | 78.57 | 47.07 | 10.20 | 3.34 | 43.75 | 57.11 | 68.63 | 76.86 |
| 4p3v(A) | 37.48 | 25.16 | 22.91 | 28.59 | 35.16 | 16.58 | 41.23 | 21.73 | 22.75 | 31.47 | 38.98 | 13.33 | 36.40 | 19.53 | 24.41 | 31.73 | 39.24 | 32.04 |
| 4p3v(A)+ENM | 4.97 | 1.48 | 65.49 | 76.47 | 84.53 | 40.45 | 7.02 | 2.59 | 49.35 | 66.32 | 78.53 | 28.35 | 10.35 | 3.38 | 43.62 | 56.99 | 68.55 | 62.44 |
| 4p3v(A)+R | 33.45 | 19.95 | 25.86 | 32.22 | 39.16 | 16.32 | 36.98 | 11.59 | 27.03 | 36.70 | 44.48 | 12.48 | 34.84 | 18.22 | 25.27 | 32.77 | 40.29 | 29.56 |
| 4p3v(A)+R+F | 35.65 | 23.47 | 23.99 | 29.97 | 36.64 | 11.10 | 38.34 | 14.47 | 25.89 | 35.19 | 42.80 | 9.35 | 36.77 | 20.76 | 23.94 | 31.23 | 38.58 | 19.75 |
| 4p3v(A)+R+F+ENM | 4.98 | 1.48 | 65.62 | 76.56 | 84.55 | 32.37 | 6.99 | 2.57 | 49.62 | 66.55 | 78.72 | 23.43 | 10.43 | 3.38 | 43.57 | 56.95 | 68.56 | 42.36 |
| 4p3v(M) | 6.02 | 1.71 | 60.86 | 72.56 | 81.73 | 35.18 | 8.36 | 2.74 | 47.15 | 64.06 | 76.59 | 25.15 | 12.09 | 3.62 | 41.78 | 55.06 | 66.67 | 54.89 |
| 4p3v(M)+ENM | 5.28 | 1.58 | 63.61 | 74.96 | 83.46 | 48.83 | 7.21 | 2.63 | 48.69 | 65.57 | 77.95 | 34.93 | 10.67 | 3.43 | 43.31 | 56.61 | 68.08 | 70.87 |
| 4p3v(M)+R | 5.49 | 1.56 | 63.66 | 74.88 | 83.37 | 41.52 | 7.69 | 2.63 | 48.69 | 65.67 | 77.99 | 30.76 | 11.17 | 3.46 | 43.28 | 56.39 | 67.88 | 60.40 |
| 4p3v(M)+R+F | 5.48 | 1.54 | 63.99 | 75.14 | 83.51 | 30.88 | 7.75 | 2.62 | 48.77 | 65.71 | 77.98 | 22.72 | 11.16 | 3.48 | 43.04 | 56.28 | 67.73 | 42.38 |
| 4p3v(M)+R+F+ENM | 5.01 | 1.50 | 65.18 | 76.16 | 84.28 | 44.51 | 6.99 | 2.58 | 49.43 | 66.31 | 78.56 | 32.48 | 10.25 | 3.39 | 43.74 | 57.21 | 68.80 | 58.08 |
| 4p3v(M) | 5.55 | 1.68 | 61.86 | 73.73 | 82.78 | 83.70 | 7.65 | 2.67 | 48.16 | 65.25 | 77.76 | 59.23 | 11.13 | 3.47 | 42.74 | 56.11 | 67.74 | 125.02 |
| 4p3v(M)+ENM | 4.99 | 1.56 | 64.02 | 75.51 | 84.01 | 125.66 | 6.88 | 2.60 | 49.11 | 66.13 | 78.51 | 89.05 | 10.26 | 3.35 | 43.66 | 56.99 | 68.55 | 175.54 |
| 4p3v(M)+R | 4.86 | 1.47 | 65.97 | 77.00 | 85.01 | 100.61 | 7.10 | 2.56 | 49.79 | 66.89 | 79.08 | 73.94 | 10.37 | 3.34 | 43.93 | 57.26 | 68.82 | 138.53 |
| 4p3v(M)+R+F | 4.92 | 1.47 | 65.85 | 76.87 | 84.90 | 71.73 | 7.19 | 2.56 | 49.71 | 66.75 | 78.95 | 52.84 | 10.52 | 3.36 | 43.79 | 57.17 | 68.73 | 92.89 |
| 4p3v(M)+R+F+ENM | 4.66 | 1.46 | 66.12 | 77.14 | 85.15 | 112.60 | 6.65 | 2.55 | 49.87 | 66.95 | 79.20 | 81.98 | 9.96 | 3.32 | 44.02 | 57.42 | 68.97 | 139.19 |
| PhotoTourism [IMC2020] | Cambridge Landmarks [kendall2015cambridge] | Aachen Day-Night v1.1 [zhang2021aachen] | ||||||||||||||||
| Estimator | AVG | MED | AUC@5 | @10 | @20 | Runtime | AVG | MED | AUC@5 | @10 | @20 | Runtime | AVG | MED | AUC@5 | @10 | @20 | Runtime |
| 4p3v(HC) [Hruby_cvpr2022] | 6.37 | 1.62 | 62.41 | 73.53 | 81.94 | 64.10 | 8.73 | 2.67 | 47.99 | 64.58 | 76.65 | 60.11 | 12.43 | 3.55 | 42.28 | 55.20 | 66.55 | 67.26 |
[…]
| 4p3v(M) | 6.02 | 1.71 | 60.86 | 72.56 | 81.73 | 35.18 | 8.36 | 2.74 | 47.15 | 64.06 | 76.59 | 25.15 | 12.09 | 3.62 | 41.78 | 55.06 | 66.67 | 54.89 |
[…]
| 4p3v(M) | 5.55 | 1.68 | 61.86 | 73.73 | 82.78 | 83.70 | 7.65 | 2.67 | 48.16 | 65.25 | 77.76 | 59.23 | 11.13 | 3.47 | 42.74 | 56.11 | 67.74 | 125.02 |
[…]
Tab. 3 shows the results for PoseLib RANSAC with early termination and 5px epipolar threshold. Results for GC-RANSAC are in SM. With the suggested modifications (see Sec. 3.2) all proposed solvers outperform the state-of-the-art 4p3v(HC) solver [Hruby_cvpr2022] in terms of pose accuracy with many of the variants also achieving faster runtimes.
Using filtering (+F) improves the run-time of RANSAC at the cost of a decrease in pose accuracy. Still, our 4p3v(M)+R+F solver outperforms 4p3v(HC) in terms of both accuracy and run-time. The 4p3v(M) and 4p3v(M)+R solvers clearly improve upon the 4p3v(M) solvers, albeit at an increased run-time. Still the 4p3v(M)+R+F solver provides, in general, the best speed-accuracy trade-off. The results also show that early non-minimal refinement (+ENM) is necessary for the 4p3v(A) solver to perform well. For the other approximate solvers, this form of refinement also leads to improved accuracy at some cost to the runtime. Assuming that the 5pt+P3P solver samples five points from which at least four have projections in all three cameras, the suggested pose refinement (+R) and filtering (+F) can also be applied to the baseline 5pt+P3P by using the fourth (potentially the fifth) point in the third view. The suggested modifications improve the performance of the 5pt+P3P solver, with the 5pt+P3P+R+F+ENM solver having an accuracy similar to that of the 4p3v(M)+R+F+ENM solver, however at slower run-times, Moreover, it is less accurate than the proposed 4p3v(M)+R+F+ENM solver.
We also investigate the speed-accuracy trade-off of the solvers by running PoseLib RANSAC [PoseLib] for a set of fixed numbers of iterations. Runtimes are reported for 1 core of a 2 GHz Intel Xeon Gold 6338 CPU. As shown in Fig. 4, the proposed 4p3v(M)+R+F solver consistently provides the best speed-accuracy trade-off both with and without ENM. 4p3v(A)+R+F+ENM provides a similar performance, typically beating 4p3v(HC) [Hruby_cvpr2022]. However, as shown in Fig. 4(d) for the St. Mary’s Church scene [kendall2015cambridge], it may perform worse for some specific scenes.
The SM provides detailed ablation studies on the significance of each individual modification (+R/+C/+ENM).
Limitations. As discussed, the accuracy of the proposed approximate solvers is scene dependent.444This weakness also applies to [Hruby_cvpr2022], since the scene needs to be similar enough to the training scenes for the MLP-based classifier to work well. However, after using the propsed ENM refitting, the scene-dependency is negligible. Especially for the 4p3v(M)+ENM solver, we have not noticed performance drops for some specific scene geometries or camera configurations. While the two-view variants of our solvers 4p(M)/4p(M) provide very similar results to the 5pt solver in two views, in this case they do not outperform the 5pt solver. In the three-view scenario, the better performance is achieved thanks to the proposed pose refinement (+R) and filtering (+F).
The proposed ENM refitting cannot be applied to the 4p3v(HC) solver [Hruby_cvpr2022], since this solver estimates the pose of all three cameras together. It is theoretically applicable to the solver proposed in [DBLP:journals/ijcv/NisterS06]. However, here ENM would have needed to be run on hundreds of candidate poses corresponding to tens-to-hundreds of sampled epipoles. This is because for one fixed point on the curve of epipoles, the error of the sampled epipole is not bounded since the true epipole can lie anywhere on the curve, even outside the image. Thus, ENM might avoid local minima present in the second step of the original solver, but would significantly slow down the solver in the first step, making it much slower than our 4p3v(M)+ENM/ 4p3v(M)+ENM solvers.
5 Conclusion
We consider the highly challenging problem of relative pose estimation of three calibrated cameras from four correspondences. We propose novel approaches that solve the problem by utilizing solvers based on approximate geometry. The best performing solver uses a simple, yet novel strategy by using mean coordinates of three input points and points in their vicinity as an approximate correspondences. Extensive experiments show that our solvers achieve state-of-the-art performance on a large variety of real scenes. At the same time, our solvers are simple to implement, especially compared to the current state-of-the-art [Hruby_cvpr2022].
Acknowledgements. This work was funded by the Czech Science Foundation (GAČR) JUNIOR STAR Grant No. 22-23183M (supporting C.T., Y.D., and Z.K.), the Grant Agency of the Czech Technical University in Prague grant no. SGS23/173/OHK3/3T/13 (supporting C.T.), the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project ”InnovAIte Slovakia, Illuminating Pathways for AI-Driven Breakthroughs” No. 09I02-03-V01-00029, the TERAIS project – a Horizon-Widera-2021 program of the European Union under the Grant agreement number 101079338 (supporting V.K. and Z.B.H.), and the EU Horizon 2020 project RICAIP grant agreement No. 857306 (supporting T.S.). Part of the research results was obtained using the computational resources procured in the national project National competence centre for high performance computing project code: 311070AKF2, funded by European Regional Development Fund, EU Structural Funds Informatization of society, Operational Program Integrated Infrastructure.
Supplementary Material
This supplementary material provides additional details and experimental results promised in the main paper: Sec. 6 provides the proof on the bound of the epipolar error of the mean point correspondence mentioned in Sec. 3.1 of the main paper, additional synthetic and noise experiments that were discussed in Sec. 4 of the main paper, accuracy-speed trade-off results for two-view approximate geometry inside RANSAC (see Sec. 4 of the main paper), and additional details and plots on more scenes for the mean point correspondence accuracy (see Sec. 4 of the main paper). Sec. 7 contains ablation studies to validate our choices regarding the modifications discussed in Sec. 3.2 of the main paper. Sec. 8 provides results for the accuracy-speed trade-off for individual scenes from the PhotoTourism [IMC2020] and Cambridge Landmarks [kendall2015cambridge] datasets, evaluations of the three-view solvers for different thresholds inside RANSAC (see Sec. 4 of the main paper), runtimes of the proposed and state-of-the-art solvers for the 4p3v problem, semi-synthetic experiments for increasing outliers ratios on a scene from PhotoTourism [IMC2020], details and results on an alternative evaluation measure (see Sec. 4 of the main paper), and results with GC-RANSAC (see Sec. 4 of the main paper). 44footnotetext: ∗ Equal contribution
6 Approximate camera geometry
In this section, in addition to the experiments presented in Sec. 4 of the main paper (paragraph “Approximate camera geometry”), we present additional experiments and results to support our idea of estimating approximate geometry in the first two views. We start with the proof on the bound of the epipolar error of the mean-point correspondence used in the proposed 4p3v(M)-based solvers (see Sec 3.1 of the main paper). Next, Sec. 6.2 discusses why the mean-point correspondence provides an additional constraint (compared to the original point correspondences) that can be used to estimate the essential matrix. Then, to further assess the accuracy of the two-view variants of our approximate solvers (outside of RANSAC), i.e., 4p(A), 4p(M), and 4p(M), we design two additional synthetic experiments (see Sec. 6.3). The goal of these synthetic experiments is to study how the accuracy of approximate solutions varies with varying properties of the scene and the cameras. Moreover, Sec. 6.3 provides a synthetic noise experiment on data extracted from a real scene from the ETH3D dataset [Schops_2017_CVPR], outside of RANSAC for the three-view solvers. Lastly, Sec. 6.4 contains a speed-accuracy evaluation of the two-view variants of the proposed solvers, inside Poselib RANSAC, and in Sec. 6.5 we study the accuracy of the mean point correspondence (as in Fig. 3 of the main paper).
6.1 Proof of the bounds on the epipolar error
While the mean point correspondence used in the 4p3v(M)-based solvers can provide a good approximation of a correct correspondence, such a correspondence can be noisy. Note that all state-of-the-art 4p3v solvers (including 4p3v(HC) [Hruby_cvpr2022] and the solver from [DBLP:journals/ijcv/NisterS06]) rely on certain approximations without establishing theoretical proofs to quantify their accuracy. In the 4p3v(HC) solver [Hruby_cvpr2022], the failures that appear quite often are usually the results of tracking a geometrically incorrect solution inside the homotopy continuation method.555The solver is tracking only one from 272 solutions of the relaxed version of the 4p3v problem and this solution does not need to be a geometrically correct one. Thus, this solution can be arbitrarily far from the correct solution. The solver from [DBLP:journals/ijcv/NisterS06] requires sampling epipoles from a -degree curve on which the true epipole must lie. For any selected point on the curve of epipoles, the error of the sampled epipole is not bounded, since the true epipole can lie anywhere on the unbounded curve. The curve is unbounded since the epipole can be located arbitrarily far away from the image center based on the relative pose of the two cameras (up to infinity for sideways motion).
In contrast, the error of our mean point correspondence is bounded. The correspondence can be seen as a correspondence of points that are projections of the mean point of three 3D points. In this case, the error of both projections and can be computed as an error that is introduced by approximating the perspective projection using the para-perspective projection. The approximation error introduced by the para-perspective projection is studied in the literature and can be found e.g. in [Zhang2014, Horaud97].
However, we can also look at the correspondence from a different point of view. We can consider this correspondence as a correspondence in which we fix a point in one view, e.g., , and generate a corresponding point in the second view. In this case, as mentioned in the main paper, it can be proven that the epipolar error of the mean point correspondence is bounded by the maximum distance of the mean point from the vertices of the triangle . Here we provide a simple proof.
Lemma 1.
Let us assume two cameras with camera centers and that observe 3D points and (see Figure 5 for an illustration). Let and be the projections of these 3D points in camera 1 and camera 2, respectively. Let be the mean point of the points and let be the essential matrix between these two cameras, i.e., a matrix that satisfies . Then the epipolar line passes through the triangle .
Proof.
The camera center and the 3D points and form a tetrahedron (see Figure 5). The projections in the first camera lie at the edges of this tetrahedron . The ray from the camera center through the mean point thus lies inside the tetrahedron and intersects the plane defined by 3D points and in a point that lies inside the triangle defined by .
The camera center and the 3D points and form a tetrahedron . Again, the projections lie at the edges of the tetrahedron . The ray passing through the camera center and the 3D point lies inside the tetrahedron and thus intersects the image plane of the second camera at a point that lies inside the triangle defined by the points . By construction, the projection of into the second camera lies on the epipolar line . Therefore, the epipolar line which is a line connecting this point and the epipole , passes through the triangle . ∎
It follows from Lemma 1 that since the epipolar line passes through the triangle , the maximum distance of the mean point to the epipolar line is equal to the maximum distance of to the vertices of the triangle.
6.2 Mean-point constraint
As already mentioned in the main paper, under the assumption of a para-perspective projection, i.e., of affine geometry, the mean point of three 3D points is projected to the mean points of the 3D points’ projections in both images [Zhang2014]. Thus, the mean point correspondence does not add a new constraint if used to estimate an affine camera. This can be easily shown. In the case of affine cameras, the essential matrix has the form
Thus the epipolar constraint for the mean point correspondence with the homogeneous coordinates ), and ):
can be written as a linear combination of the epipolar constraint for the three input points, i.e. .
This is not the case for perspective cameras. For perspective cameras, i.e., when estimating the full essential matrix , the mean point correspondence introduces an additional constraint. In this case the epipolar constraint
after expansion, contains terms . Thus, the epipolar constraint (3) is not a linear combination of the individual epipolar constraints . Therefore, the mean point correspondence provides an independent constraint when used to estimate the epipolar geometry of perspective cameras.
6.3 Synthetic Experiments
The error of the relative poses estimated with the proposed approximate 4p3v(M)-based and 4p3v(A)-based solvers depends on many aspects, e.g., the baseline and the view angles of the cameras w.r.t. the three points used to compute the mean point correspondence, the depth of these points, the size and shape of the triangles defined by the three points, the type of motion of the cameras, the depth of the scene and the distance of cameras from the scene, the level of noise in the correspondences, etc. Isolating the impact of the individual aspects, e.g., through experiments on synthetic data, is highly non-trivial (e.g., how to generate realistic synthetic scenarios that allow conclusions to generalize to real-world scenarios) and analysing the co-dependencies between different aspects on the overall performance seems to need a paper on its own. Moreover, the effect of approximation introduced by using para-perspective projection was already studied in the literature [Zhang2014, Horaud97].
In the main paper, we thus presented mostly results on real-world scenes, without trying to isolate individual factors (see Figure 3 and Table 1 in the main paper). However, we also tested interesting camera and scene setups using synthetically generated data.
We extend the synthetic experiment for increasing angles between the projection rays of the cameras in the main paper (Fig. 2 of the main paper), by investigating the impact of increased noise, alternative noise models, and higher outlier ratio. Similar to the experiment of increasing angles, we evaluate the two-view solvers (outside of RANSAC) in two additional interesting scenarios. The goal is to study the effect of the proposed approximations on the relative pose estimation under varying properties of the scene and the cameras.
Additionally, we test the performance (outside of RANSAC) of our proposed approximate solvers and the state-of-the-art solvers for the 4p3v problem w.r.t. increasing image noise added to ground-truth correspondences extracted from a scene from the ETH3D dataset [Schops_2017_CVPR]. As an evaluation metric for the two-view geometry, we use the pose error measured as [IMC2020].
Increasing angle between principal axes of cameras. In Fig. 6, we present results analogous to Fig. 2 of the main paper, but for px and px noise levels. We also include results for the uniform noise model (noise is evenly distributed in range ). As in Fig. 2 of the main paper, the synthetic data contain 20% outliers. As in the results presented in Fig. 2 of the main paper, the accuracy of both approximate solvers decreases as the angle increases, where 4p(A) demonstrates notably lower accuracy than 4p(M) and 4p(M). However, ENM singnificantly improves the accuracy, with 4p(M)+ENM achieving the same or slightly better accuracy than 5pt+ENM for angle . In Fig. 7 we present results for Gaussian noise and higher outlier ratio, in particular 40% outliers. The results are consistent with those of Fig. 6, though the increased outlier ratio leads to a higher error when using ENM.
Increasing distance of cameras to the 3D scene. It is known that the quality of the affine approximation, i.e., the approximation of the perspective projection using the para-perspective projection, depends on the distance of the points from the camera [Zhang2014]. Thus in the first experiment, we evaluate the performance of all solvers w.r.t. increasing distance of the cameras to the 3D scene.
We perform this experiment on 10k synthetically generated instances. For each of the 10k instances, we uniformly sample 3D points inside a 2000x2000x100 unit cube, and the camera centers are placed at random points with the same distance from the scene. The distances tested (in units) are . The cameras are generated such that they look towards the scene. We add 1px noise to the projected points. Note that the scene is generated such that the projections of the points cover a large portion of the image for cameras at all distances.
Fig. 8(a) shows the results of this experiment represented by the boxplot function, which shows values between the and quantiles as a box with a horizontal line at the median. As expected, the errors of the 4p(A) solver decrease as the distance of the cameras from the scene increases, since affine geometry can be better satisfied with larger distances from the scene. Without considering ENM, 5p(E) is the best performing solver. The errors of the 5p(E) solver are increasing with increasing distance of the cameras from the scene (due to the fact that fixed image noise is generating larger errors for points that are farther from cameras). This effect is less visible for the proposed 4p(M) and 4p(M) solvers since for these solvers the error is originally more dominated by the error in the mean point correspondence. When considering ENM, 5p(E), 4p(M), and 4p(M) solvers perform similarly, with 4p(M) being the most accurate for distances . The 4p(A) solver is also greatly improved when using ENM, reaching similar or even better accuracy than 5p(E) (w/o ENM), for distances .
Increasing depth of the 3D scene. In Fig. 8(b), instead of increasing the distance of the cameras to the 3D scene, we place the cameras randomly at distances between 1000 and 1200 units away from the scene, looking towards the scene, while changing the depth of the scene. In particular, the 3D points are generated uniformly at random inside a 2000x2000xdepth unit cube, where the depth of the scene is specified by the values on the x-axis. The tested depths are . We add 1px noise to the projected points. Without using ENM, the pose errors of 4p(A) are visibly decreasing as the depth of the scenes increases. This is to be expected since increasing the scene depth increases the chance of sampling four points that are more consistent with the para-perspective / affine camera model (since the points are more likely to be farther away from the cameras). The remaining solvers, that is, 5p(E), 4p(M), and 4p(M), are not significantly affected by the changes in the depth of the scene. When using ENM, all tested solvers are improved significantly, with 4p(M) being the most accurate in terms of pose error. When using ENM, the errors in the estimated poses increase with increasing scene depths, which is particularly visible for the 4p(A) solver. This behavior is due to the fact that for points farther away from the camera, the same amount of image noise (1px) has a larger impact, thus leading to the non-minimal samples being more affected by noise. Still, using ENM clearly leads to significantly smaller errors for all solvers.
Noise experiments. We test the performance of our solvers and the state-of-the-art solvers w.r.t. increasing image noise. We used the SfM model of the botanical garden scene (randomly selected from all scenes) from the ETH3D dataset [Schops_2017_CVPR] to obtain instances of 5 points in three views by identifying images in the scene that share 3D points. Perfect noise-free image correspondences are generated by projecting the 3D points into the images. We then add increasing amounts of normally distributed noise to these correspondences. We generated more than 1k instances. Note that the 4p3v(HC) solver was trained on the ETH3D dataset [Schops_2017_CVPR].
The results for increasing noise in the image points are shown in Fig. 9. The figure shows boxplots of the pose errors measured in the same way as in our experiments in the main paper (cf. Sec. 4 in the main paper), i.e., as .666Here is the error of the estimated relative rotation between cameras and , computed as the angle in the axis-angle representation of , and is the error of the estimated translation computed as the angle between the two unit vectors corresponding to the translations [IMC2020]. The errors are zoomed into an interesting interval and are shown as functions of varying Gaussian noise from px to px.
Due to the approximations used in our proposed 4p3v(M)-based and 4p3v(A)-based solvers, these solvers exhibit non-zero errors for zero noise. However, at noise levels px, our 4p3v(M)+R solvers return comparable or even better (w/ ENM) results than the 5pt+P3P solver. For noise px, also the 4p3v(M)+R solver with ENM returns slightly more accurate poses than the 5pt+P3P solver with ENM. In general, the effect of increasing image noise is less visible for approximate 4p3v(M)-based and 4p3v(A)-based solvers. In this case, the error of the approximation is dominating the error introduced by the noise in the image correspondences. While for 4p3v(A)-based solvers the approximation error is dominant at all noise levels, for 4p3v(M)-based solvers, at noise px, the error introduced by the approximate mean point correspondence (and points in their vicinity in -based solvers) is suppressed by the error introduced by noise in the remaining point correspondences.777This can be seen from the comparable pose accuracy of the 5pt+P3P and 4p3v(M) solvers for px noise, and the comparable accuracy of the 5pt+P3P and 4p3v(M) solvers for px noise. Note that, although the pose errors for the 4p3V(A)+R+ENM solver are higher than those of the rest of the solvers, as shown in our real experiments, this solver still returns reasonably low errors to provide local optimization (LO) within RANSAC with a good initialization in real-world settings. Further, note that the 5pt+P3P solver samples one more point (real correspondence) in the first two cameras, and these points are affected only by the considered noise. This shows that the mean point correspondence used in the 4p3v(M)-based solver is a good approximation to a real correspondence. The recent state-of-the-art 4p3v(HC) solver [Hruby_cvpr2022] is failing in about 50% of the instances for noiseless data, even though the solver was trained on the ETH3D dataset. Thus, the median errors are significantly larger than the median errors of the remaining solvers.
6.4 Two-view approximate solutions inside RANSAC
In the next experiment, we evaluate the discussed two-view solvers, i.e., the 5p(E), 4p(M), 4p(M), and 4p(A) solvers, inside RANSAC as well. This experiment indicates how the proposed approximate solvers would have behaved if used as two-view solvers. Note that in the two-view case, the proposed filtering (+F) and refinement (+R) using the point correspondence in the third view are not applicable.
Fig. 10 shows the speed-accuracy evaluation of the solvers for the problem of two view relative pose estimation evaluated using PoseLib RANSAC [PoseLib] on 12 scenes from the Phototourism dataset [IMC2020] (excluding St. Peter’s Square) with pairwise point correspondences obtained using [detone2018superpoint, lindenberger2023lightglue]. The statistic reported is the AUC@10∘ of the pose error for a varied number of RANSAC iterations () and a fixed epipolar threshold of 5px. The upper figure is a zoom-in of the lower figure to an interesting interval, where differences between the solvers are more visible. Although all proposed solvers (except the 4p(A) solver without ENM) have a performance comparable to that of the state-of-the-art two view 5pt solver, the 5pt solver is the best performing one for the two-view scenario. This result is not surprising, given the well-known good performance (in terms of speed and accuracy) of the 5pt solver and not very high outlier contamination of the data (for which sampling one point less would have potentially had a more visible effect). It also indicates that the proposed modifications, i.e., the filtering +F the and refinement +R, for the three-view scenario are important and are making the proposed 4p3v approximate solvers practical and more precise than the 5pt+P3P sovler.
6.5 Accuracy of the mean point correspondence
Fig. 3 in the main paper showed results obtained by establishing correspondences between the mean of the triangle in one image and various points in the triangle in the second image. We expressed points in the second triangle via their barycentric coordinates and uniformly sample barycentric coordinates , such that (ensuring points inside the triangle). The 3rd coordinate is given as . For each correspondence, we measured the symmetric epipolar error w.r.t. the ground truth pose, translation and rotation errors, and the percentage of inliers. Fig. 3 in the main paper showed the rotation error and percentage of inliers, as observed for the St. Peter’s Square scene from the PhotoTourism dataset [IMC2020]. Here, Fig. 11 shows the same statistics, including translation and symmetric epipolar errors, for the St. Peter’s Square scene already used in the main paper (Fig. 11 (top row)), and two more scenes from the PhotoTourism dataset: Sacre Coeur (Fig. 11 (middle row)), and Temple Nara Japan (Fig. 11 (bottom row)).
As with Fig. 3 in the main paper, to suppress the effect of discrete sampling, for each metric, we fit a 2D Gaussian distribution and report the mean value (in barycentric coordinates) as numbers in brackets in the caption of the figure. As can be seen, the same conclusion can be drawn from Fig. 11 as from Fig. 3 in the main paper: The optima of the studied metrics are reached very close to the mean point of the triangles, which has barycentric coordinates . This validates our approach of using the mean point correspondence as an approximate correspondence in our 4p3v(M)-based solvers.
| Estimator | AVG | MED | AUC@5 | @10 | @20 | |
| 4p3v(M) | 0.2 | 4.03 | 2.09 | 57.81 | 73.48 | 84.67 |
| 0.1 | 3.99 | 2.03 | 58.29 | 73.71 | 84.78 | |
| 0.09 | 3.99 | 2.02 | 58.52 | 73.95 | 84.93 | |
| 0.08 | 3.95 | 2.04 | 58.66 | 74.11 | 85.04 | |
| 0.07 | 4.02 | 2.03 | 58.63 | 74.01 | 84.98 | |
| 0.06 | 4.01 | 2.01 | 58.62 | 73.92 | 84.90 | |
| 0.05 | 3.98 | 1.98 | 58.94 | 74.19 | 85.04 | |
| 0.01 | 4.01 | 2.07 | 57.90 | 73.60 | 84.75 | |
| 0.005 | 4.12 | 2.07 | 57.54 | 73.23 | 84.55 | |
| 0.001 | 4.28 | 2.14 | 56.65 | 72.51 | 84.08 | |
| 4p3v(M)+R | 0.2 | 3.75 | 1.87 | 61.36 | 75.83 | 86.07 |
| 0.1 | 3.75 | 1.87 | 61.33 | 75.78 | 86.03 | |
| 0.09 | 3.74 | 1.87 | 61.39 | 75.81 | 86.02 | |
| 0.08 | 3.71 | 1.87 | 61.45 | 75.90 | 86.16 | |
| 0.07 | 3.76 | 1.87 | 61.41 | 75.82 | 86.06 | |
| 0.06 | 3.79 | 1.87 | 61.40 | 75.81 | 86.04 | |
| 0.05 | 3.75 | 1.86 | 61.38 | 75.79 | 86.03 | |
| 0.01 | 3.75 | 1.87 | 61.25 | 75.70 | 85.95 | |
| 0.005 | 3.73 | 1.87 | 61.22 | 75.75 | 86.03 | |
| 0.001 | 3.86 | 1.90 | 60.82 | 75.46 | 85.80 | |
| 4p3v(M)+R+F | 0.2 | 3.78 | 1.88 | 61.12 | 75.70 | 85.96 |
| 0.1 | 3.77 | 1.87 | 61.16 | 75.68 | 85.95 | |
| 0.09 | 3.76 | 1.87 | 61.24 | 75.70 | 85.92 | |
| 0.08 | 3.73 | 1.87 | 61.30 | 75.78 | 86.07 | |
| 0.07 | 3.77 | 1.87 | 61.30 | 75.71 | 85.99 | |
| 0.06 | 3.80 | 1.87 | 61.27 | 75.75 | 85.99 | |
| 0.05 | 3.75 | 1.87 | 61.28 | 75.73 | 85.98 | |
| 0.01 | 3.78 | 1.88 | 61.11 | 75.62 | 85.89 | |
| 0.005 | 3.76 | 1.89 | 61.09 | 75.65 | 85.95 | |
| 0.001 | 3.90 | 1.91 | 60.53 | 75.26 | 85.67 |
7 Ablation studies
This section contains ablation studies to validate our choices in modifications discussed in Sec. 3.2 of the main paper.
Validation of . We tested our -based solvers for different values of and measured their performance. In general, there is no common value of the shift that leads to the best results on all datasets. This is expected since the precision of the mean-point correspondence depends on many different factors, e.g., the viewing angles of the cameras, the type of the motion, the depth and spatial distributions of the 3D points, etc. We set the value for by evaluating their effects on the St. Peter’s Square scene from the PhotoTourism dataset [IMC2020], which we used for validation only and did not include it in the other results for PhotoTourism [IMC2020] in the paper. Tab. 4 shows how the different settings of the scale of the shift affect the accuracy of the -based solvers. Based on these experiments, we use as it typically provides the best or the second best results for all variants of the 4p3v(M)-based solvers. However, note that 4p3v(M)-based solvers achieve a very similar accuracy even with different settings of . Thus, we can conclude that the choice of is not critical. In some scenarios, the choice of the optimal parameter may be more scene-dependent and could potentially be set using learning-based approaches.
Refinement validation. We also perform validation of the total number of LM steps in the refinement (+R). Again, we used the St. Peter’s Square scene from the PhotoTourism dataset [IMC2020] for validation. The results of this experiment are shown in Fig. 12. We chose the value of 2 for other experiments as it provides good speed-accuracy trade-off across a range of RANSAC iterations. However, we note that other settings provide very similar performance.
Validation of +F/+R/+ENM. Fig. 13 ablates the impact of individual modifications (+R/+C/+ENM) proposed in Sec. 3.2 in the main paper on the speed-accuracy trade-off. It especially highlights the importance of the refinement using the 4th point in the third view (+R). Fig. 13 also shows the performance of the top-performing solvers when different maximum epipolar thresholds are used within RANSAC. Compared to 4p3v(A)+R+F+ENM, the proposed 4p3v(M)-based solvers are not as sensitive to the selection of the epipolar threshold. The results presented in Fig. 13 were obtained on the PhotoTourism dataset. Fig. 14 shows results of the same ablation study, focused on the 4p3v(M)-based solvers, on the Cambridge Landmarks and Aachen Day-Night v1.1 datasets.
8 Experiments on real data
In this section, we aim to further study the performance of the proposed methods, supplementing Sec. 4 (paragraph “Experiments on real data”) of main paper with more detailed evaluations. Section 8.1 presents results on individual scenes, extending the analysis in Fig. 4 of the main paper. Section 8.2 investigates the impact of varying the RANSAC epipolar threshold on solver performance. These experiments extend Fig. 13 (bottom) by comparing additional solvers across all three datasets. Section 8.3 evaluates and compares the run-times of each of the proposed and state-of-the-art solvers. Section 8.4 explores the robustness of the solvers under varying inlier ratios using semi-synthetic data. In Section 8.5, we provide results using Poselib RANSAC [PoseLib] for an alternative pose error metric that considers errors across all three camera pairs. Section 8.7 evaluates the solvers within the GC-RANSAC [barath2017graph] framework for all three datasets.
8.1 Results on individual scenes
Fig. 4 in the main paper showed results jointly on all PhotoTourism [IMC2020] scenes (except St. Peter’s Square), jointly on the 5 Cambridge Landmarks [kendall2015cambridge] scenes (except the Street scene, which is commonly not used due to issues with its ground truth), and Aachen Day-Night v1.1 [zhang2021aachen]. It also showed results on one individual scene from [kendall2015cambridge], i.e., the St. Mary’s Church scene. In Fig. 15, we provide results for the accuracy-speed trade-off evaluation for all remaining individual scenes of PhotoTourism [IMC2020] and Cambridge Landmarks [kendall2015cambridge].
As discussed in the main paper, the accuracy of the proposed approximate solvers is scene-dependent. This also applies to the state-of-the-art 4p3v(HC) solver [Hruby_cvpr2022], since in this solver the scene needs to be similar enough to the training scenes for the MLP-based classifier to work well. The proposed ENM refitting suppresses to some extent the scene dependency of the proposed 4p3v(M)-based and 4p3v(A)-based solvers. It can be seen that the proposed 4p3v(M)+R+F solver consistently provides the best speed-accuracy trade-off both with and without ENM accross all scenes. 4p3v(A)+R+F+ENM provides a similar performance, typically beating 4p3v(HC) [Hruby_cvpr2022]. However, it may perform worse for some specific scenes, e.g., Shop Facade and Palace of Westminster. In general, the results on individual scenes are consistent with the results from the main paper.
| Phototourism [IMC2020] | ![]() |
| Cambridge [kendall2015cambridge] | ![]() |
| Aachen [zhang2021aachen] | ![]() |
8.2 RANSAC threshold sensitivity
In Fig. 13 (bottom), we provided experiments showing how the performance of the selected methods changes when we vary the RANSAC epipolar threshold. In Fig. 16 we provide more extensive results comparing the methods with different thresholds on all three datsets.
Similar to the results presented in the main paper, 4p3v(M)+R+F in both variants (with and without ENM) shows consistently good performance even when using a different threshold in RANSAC. In contrast, 4p3v(A)+R+F+ENM performs worse than 4p3v(M)+R+F when considering a higher epipolar threshold in RANSAC.
8.3 Solver run-times
In this section, we present run-times of the proposed solvers as well as the state-of-the-art solvers for the relative pose problem of three calibrated cameras. While the main paper reports run-time results for full RANSAC-based estimation, we now report the run-times of the individual solvers outside of RANSAC. To measure the run-times of the solvers888Note that for 4p3v(HC) solver, in Tab 5, the time needed to load the weights of the network (or any other required data) is not added to the runtime of the solver. The data are loaded once per RANSAC, and the loading takes on average 45ms., we calculated the average run-time of each solver on more than 50k instances of the Sacre Coeur scene of the PhotoTourism dataset [snavely2006photo]. The run-times are reported in Table 5. The experiments were performed on an Intel(R) Core(TM) i9-10900X CPU @ 3.70GHz. In general, the implementations of all proposed solvers are not optimized for speed, and we still see room for speeding them up.
| 5pt+P3P | 4p3v(HC) | 4p3v(M) | 4p3v(M) | 4p3v(A) | |
| Time (s) | 77.90 | 66.06 | 83.92 | 218.71 | 61.12 |
8.4 Outlier experiments
To show how the different solvers perform even under varying inlier ratios, we perform a semi-synthetic experiment. We use the Notre Dame scene from PhotoTourism [IMC2020]. We keep all inlier triplets w.r.t. a 5px epipolar threshold using the ground truth poses. We add additional synthetic outlier correspondences by generating random points in all three views. This allows us to study how the different methods perform when the inlier ratio changes. The results are shown in Fig. 17. As expected, the performance of all solvers decreases with lower inlier ratios. We also observe that 4p3v(M)+R+F performs well even with a low inlier ratio. In contrast, the relative performance of 4p3v(A)+R+F+ENM worsens with a decreased inlier ratio. However, we note that even with an inlier ratio of 40%, it still results in performance comparable to the baseline 5pt+P3P solver. This suggests that a high inlier ratio is not necessary for the ENM to work well in conjunction with the solver 4p3v(A).
| Phototourism [IMC2020] | Cambridge Landmarks [kendall2015cambridge] | Aachen Day-Night v1.1 [zhang2021aachen] | ||||||||||||||||
| Estimator | AVG | MED | AUC@5 | @10 | @20 | Runtime (ms) | AVG | MED | AUC@5 | @10 | @20 | Runtime (ms) | AVG | MED | AUC@5 | @10 | @20 | Runtime (ms) |
| 4p3v(HC) [Hruby_cvpr2022] | 10.28 | 2.81 | 46.97 | 61.37 | 73.16 | 64.10 | 13.95 | 4.58 | 29.54 | 48.68 | 65.12 | 60.11 | 23.46 | 6.58 | 28.56 | 41.20 | 53.12 | 67.26 |
| 5pt+P3P | 9.02 | 2.81 | 47.06 | 61.89 | 74.12 | 33.77 | 12.39 | 4.57 | 29.56 | 49.03 | 65.94 | 24.04 | 21.17 | 6.33 | 29.07 | 41.90 | 54.13 | 53.34 |
| 5pt+P3P+ENM | 8.77 | 2.73 | 47.93 | 62.68 | 74.73 | 48.79 | 12.00 | 4.52 | 29.76 | 49.35 | 66.25 | 34.82 | 21.39 | 6.42 | 28.91 | 41.73 | 53.78 | 71.61 |
| 4p3v(A) | 58.51 | 46.32 | 16.16 | 21.93 | 27.71 | 16.58 | 59.75 | 43.50 | 13.59 | 22.93 | 31.44 | 13.33 | 53.53 | 41.38 | 17.05 | 23.62 | 30.10 | 32.04 |
| 4p3v(A)+ENM | 8.44 | 2.60 | 49.46 | 64.10 | 75.85 | 40.45 | 11.86 | 4.46 | 30.17 | 49.89 | 66.74 | 28.35 | 21.00 | 6.23 | 29.08 | 42.05 | 54.32 | 62.44 |
| 4p3v(A)+R | 53.59 | 38.24 | 18.31 | 24.91 | 31.34 | 16.32 | 54.17 | 24.26 | 16.60 | 27.32 | 36.63 | 12.48 | 51.81 | 38.69 | 17.55 | 24.35 | 31.05 | 29.56 |
| 4p3v(A)+R+F | 56.31 | 43.55 | 16.89 | 23.07 | 29.15 | 11.10 | 55.99 | 30.64 | 15.71 | 26.07 | 35.10 | 9.35 | 54.12 | 42.69 | 16.64 | 23.04 | 29.50 | 19.75 |
| 4p3v(A)+R+F+ENM | 8.45 | 2.59 | 49.59 | 64.22 | 75.92 | 32.37 | 11.78 | 4.41 | 30.41 | 50.12 | 66.93 | 23.43 | 21.11 | 6.29 | 29.02 | 41.97 | 54.18 | 42.36 |
| 4p3v(M) | 9.98 | 3.01 | 45.16 | 60.02 | 72.55 | 35.18 | 13.63 | 4.73 | 28.66 | 47.83 | 64.70 | 25.15 | 23.44 | 6.82 | 27.90 | 40.53 | 52.69 | 54.89 |
| 4p3v(M)+ENM | 8.86 | 2.75 | 47.68 | 62.54 | 74.67 | 48.83 | 12.10 | 4.52 | 29.80 | 49.31 | 66.17 | 34.93 | 21.51 | 6.40 | 28.91 | 41.73 | 53.83 | 70.87 |
| 4p3v(M)+R | 9.24 | 2.74 | 47.74 | 62.44 | 74.47 | 41.52 | 12.76 | 4.53 | 29.77 | 49.27 | 66.12 | 30.76 | 22.06 | 6.39 | 28.92 | 41.69 | 53.70 | 60.40 |
| 4p3v(M)+R+F | 9.19 | 2.71 | 48.10 | 62.73 | 74.69 | 30.88 | 12.84 | 4.53 | 29.74 | 49.29 | 66.11 | 22.72 | 22.06 | 6.42 | 28.70 | 41.53 | 53.62 | 42.38 |
| 4p3v(M)+R+F+ENM | 8.52 | 2.62 | 49.16 | 63.82 | 75.60 | 44.51 | 11.79 | 4.44 | 30.27 | 49.91 | 66.72 | 32.48 | 20.93 | 6.21 | 29.16 | 42.16 | 54.46 | 58.08 |
| 4p3v(M) | 9.28 | 2.94 | 45.92 | 61.05 | 73.66 | 83.70 | 12.68 | 4.59 | 29.35 | 48.88 | 65.89 | 59.23 | 22.09 | 6.42 | 28.60 | 41.55 | 53.76 | 125.02 |
| 4p3v(M)+ENM | 8.44 | 2.73 | 48.01 | 63.02 | 75.20 | 125.66 | 11.61 | 4.47 | 30.04 | 49.75 | 66.68 | 89.05 | 20.80 | 6.22 | 29.26 | 42.16 | 54.32 | 175.54 |
| 4p3v(M)+R | 8.32 | 2.58 | 49.79 | 64.53 | 76.30 | 100.61 | 11.93 | 4.40 | 30.47 | 50.35 | 67.22 | 73.94 | 21.15 | 6.12 | 29.35 | 42.31 | 54.53 | 138.53 |
| 4p3v(M)+R+F | 8.39 | 2.58 | 49.69 | 64.41 | 76.18 | 71.73 | 12.06 | 4.41 | 30.42 | 50.25 | 67.08 | 52.84 | 21.38 | 6.15 | 29.24 | 42.22 | 54.43 | 92.89 |
| 4p3v(M)+R+F+ENM | 7.99 | 2.56 | 49.93 | 64.67 | 76.45 | 112.60 | 11.36 | 4.39 | 30.54 | 50.40 | 67.30 | 81.98 | 20.64 | 6.08 | 29.40 | 42.48 | 54.67 | 139.19 |
8.5 Alternative evaluation measure
For the evaluation in the main paper, we defined the pose error as , where and are the angular errors of rotation and translation (both in degrees) for camera pair . The 4p3v problem also includes the estimation of and since the relative scale of and is recovered. We therefore also present results for the pose error defined as
The results equivalent to Tab. 2 from the main paper using this pose error definition are presented in Tab. 6. A speed-accuracy comparison equivalent to Fig. 4 in the main paper is presented in Fig. 18. The overall ranking of the methods remains the same under both the metric used in the main paper and the alternative described in this section.
8.6 Comparison with [DBLP:journals/ijcv/NisterS06]
The authors of [DBLP:journals/ijcv/NisterS06] kindly shared their source code with us. Unfortunately, we were not able to run the part of the code that samples epipole candidates from a 10-degree polynomial curve (appropriately sampling the curve is hard as the epipole can be arbitrarily far from the image center). At the same time, the authors were also not able to run it.
Based on the working parts of the code, we tested an oracle version of [DBLP:journals/ijcv/NisterS06], where instead of sampling the 10-degree polynomial curve, the oracle gives us the correct epipole. Given a sample close to the correct epipole, [DBLP:journals/ijcv/NisterS06] performs comparable to our M-based solvers. In practice it is hard to find good samples (the epipole can be arbitrarily far from the image center). [DBLP:journals/ijcv/NisterS06] report using 40-1,000 samples with additional local optimization for robust estimation. Even then, [DBLP:journals/ijcv/NisterS06] show that this approach performs worse than 5pt+P3P on synthetic data. In contrast, two additional samples in our -based solvers already lead to better accuracy than 5pt+P3P.
| Phototourism [IMC2020] | ||||||
| Estimator | AVG | MED | AUC@5 | @10 | @20 | Runtime (s) |
| 4p3v(HC) [Hruby_cvpr2022] | 5.34 | 1.89 | 59.19 | 72.25 | 82.10 | 2.95 |
| 5pt+P3P | 5.18 | 1.86 | 59.48 | 72.35 | 82.16 | 1.99 |
| 5pt+P3P+ENM | 5.15 | 1.86 | 59.53 | 72.39 | 82.20 | 2.05 |
| 4p3v(A) | 5.75 | 1.90 | 59.02 | 71.87 | 81.61 | 2.34 |
| 4p3v(A)+R | 5.69 | 1.72 | 61.59 | 73.83 | 82.88 | 2.98 |
| 4p3v(A)+R+F | 5.76 | 1.73 | 61.53 | 73.79 | 82.83 | 3.00 |
| 4p3v(A)+R+F+ENM | 5.34 | 1.72 | 61.71 | 74.00 | 83.09 | 2.87 |
| 4p3v(M) | 5.21 | 1.88 | 59.35 | 72.29 | 82.15 | 1.99 |
| 4p3v(M)+R | 4.94 | 1.72 | 61.91 | 74.21 | 83.36 | 2.71 |
| 4p3v(M)+R+F | 4.94 | 1.72 | 61.88 | 74.22 | 83.35 | 2.71 |
| 4p3v(M)+R+F+ENM | 4.93 | 1.72 | 61.84 | 74.21 | 83.38 | 2.73 |
| 4p3v(M) | 5.09 | 1.89 | 59.41 | 72.50 | 82.38 | 2.01 |
| 4p3v(M)+R | 4.90 | 1.71 | 61.90 | 74.26 | 83.42 | 2.76 |
| 4p3v(M)+R+F | 4.88 | 1.71 | 61.95 | 74.31 | 83.47 | 2.75 |
| 4p3v(M)+R+F+ENM | 4.86 | 1.72 | 61.90 | 74.29 | 83.48 | 2.84 |
| Cambridge Landmarks [kendall2015cambridge] | ||||||
| 4p3v(HC) [Hruby_cvpr2022] | 8.13 | 3.05 | 43.75 | 60.73 | 73.93 | 2.37 |
| 5pt+P3P | 8.01 | 3.09 | 43.17 | 60.17 | 73.67 | 2.31 |
| 5pt+P3P+ENM | 8.09 | 3.11 | 43.15 | 60.02 | 73.50 | 2.48 |
| 4p3v(A) | 8.59 | 3.11 | 43.16 | 59.98 | 73.15 | 2.62 |
| 4p3v(A)+R | 7.95 | 2.80 | 46.27 | 63.20 | 75.86 | 2.96 |
| 4p3v(A)+R+F | 8.05 | 2.81 | 46.28 | 63.17 | 75.75 | 2.98 |
| 4p3v(A)+R+F+ENM | 7.75 | 2.81 | 46.38 | 63.20 | 75.86 | 2.98 |
| 4p3v(M) | 7.95 | 3.08 | 43.32 | 60.37 | 73.75 | 2.34 |
| 4p3v(M)+R | 7.22 | 2.80 | 46.48 | 63.52 | 76.30 | 2.86 |
| 4p3v(M)+R+F | 8.05 | 2.81 | 46.28 | 63.17 | 75.75 | 2.98 |
| 4p3v(M)+R+F+ENM | 7.75 | 2.81 | 46.38 | 63.20 | 75.86 | 2.98 |
| 4p3v(M) | 7.82 | 3.06 | 43.58 | 60.60 | 73.99 | 2.41 |
| 4p3v(M)+R | 7.16 | 2.80 | 46.52 | 63.51 | 76.25 | 3.01 |
| 4p3v(M)+R+F | 7.12 | 2.81 | 46.33 | 63.47 | 76.35 | 2.95 |
| 4p3v(M)+R+F+ENM | 7.19 | 2.80 | 46.42 | 63.44 | 76.24 | 3.29 |
| Aachen Day-Night v1.1 [zhang2021aachen] | ||||||
| 4p3v(HC) [Hruby_cvpr2022] | 10.73 | 3.84 | 39.94 | 53.01 | 64.77 | 1.90 |
| 5pt+P3P | 10.76 | 3.91 | 39.54 | 52.67 | 64.56 | 1.77 |
| 5pt+P3P+ENM | 10.78 | 3.79 | 39.86 | 53.06 | 64.79 | 1.89 |
| 4p3v(A) | 11.06 | 3.88 | 39.54 | 52.63 | 64.19 | 2.24 |
| 4p3v(A)+R | 10.09 | 3.50 | 42.64 | 55.75 | 67.01 | 3.32 |
| 4p3v(A)+R+F | 10.22 | 3.49 | 42.58 | 55.75 | 66.87 | 3.38 |
| 4p3v(A)+R+F+ENM | 10.23 | 3.48 | 42.58 | 55.75 | 66.88 | 3.32 |
| 4p3v(M) | 10.62 | 3.83 | 39.62 | 52.92 | 64.74 | 1.87 |
| 4p3v(M)+R | 10.11 | 3.46 | 42.69 | 55.90 | 67.11 | 3.17 |
| 4p3v(M)+R+F | 10.22 | 3.53 | 42.58 | 55.62 | 66.71 | 3.18 |
| 4p3v(M)+R+F+ENM | 10.06 | 3.52 | 42.55 | 55.76 | 67.00 | 3.20 |
| 4p3v(M) | 10.55 | 3.84 | 39.86 | 53.30 | 65.15 | 1.89 |
| 4p3v(M)+R | 10.04 | 3.45 | 42.91 | 56.02 | 67.15 | 3.24 |
| 4p3v(M)+R+F | 10.15 | 3.47 | 42.65 | 55.76 | 66.94 | 3.17 |
| 4p3v(M)+R+F+ENM | 10.04 | 3.50 | 42.67 | 55.82 | 66.95 | 3.33 |
8.7 GC-RANSAC
Besides PoseLib’s RANSAC implementation, we also evaluated and compared our proposed solvers with the state-of-the-art solvers inside the GC-RANSAC [barath2017graph] framework.
In GC-RANSAC, local optimization (LO) is performed using non-minimal solvers that fit models to larger-than-minimal samples. We use the non-minimal version999The non-minimal version of the 5pt solver [Nister-5pt-PAMI-2004] uses the last four vectors from the SVD/QR decomposition of a matrix instead of the 4-dim null space of a matrix to parameterize the unknown essential matrix. of the 5pt solver [Nister-5pt-PAMI-2004] and the non-minimal absolute pose DLSPnP [dlspnp] solver.101010This may not be the most efficient way how to perform non-minimal refitting. However, since all methods use the same LO, it is sufficient for a fair comparison. In contrast to LO used in Poselib RANSAC, where the estimated model is used as an initialization of the Levenberg–Marquardt algorithm, in GC-RANSAC, the estimated model is used only to score inliers.
Tab. 7 shows the results for GC-RANSAC with PROSAC sampling [prosac2005chum] and a 5px epipolar threshold for all scenes from the PhotoTourism [IMC2020] dataset, 5 scenes from the Cambridge Landmarks [kendall2015cambridge] dataset, and the Aachen Day-Night v1.1 [zhang2021aachen] dataset. Similarly to what was observed for Poselib RANSAC (see Table 2 in the main paper), with the suggested modifications, all the proposed solvers outperform the state-of-the-art 4p3v(HC) solver [Hruby_cvpr2022] and the baseline 5pt+P3P solver in terms of pose accuracy with comparable runtimes. Again, the -based solvers provide, in general, the best speed-accuracy trade-off.
Due to a different LO, there are several differences compared to the results from Poselib RANSAC. Since in GC-RANSAC the estimated model is used only to score inliers, it does not need to be as precise as in Poselib RANSAC. Thus, even 4p3v(A) solver without any modification provides reasonably precise results.111111Note that the model of 4p3v(A) is refitted in the LO step with the non-minimal 5pt solver. This is similar to the refitting used in ENM, i.e., GC-RANSAC’s local optimization includes some form of ENM, explaining why the 4p3v(A) performs quite well. In contrast to this, in Poselib RANSAC, the 4p3v(A) solver without any modification results in large errors (see Table 2 in the main paper). Without refitting using ENM, the affine model estimated for the first two views in the 4p3v(A) solver is not sufficiently precise to provide a good initialization for Levenberg–Marquardt-based optimization in Poselib’s LO. Still, even for GC-RANSAC, the pure 4p3v(A) solver performs worse than the remaining variants of the proposed 4p3v(A)-based and 4p3v(M)-based solvers.
Another difference is in refitting using ENM. For GC-RANSAC, the effect of ENM is not as significant as for Poselib RANSAC. When applied without refinement (+R), the early non-minimal refitting (ENM), in general, increases the precision of solvers. When combined with +R, the improvement is not very visible. The reason is that the model returned after refining the initial, approximate model estimated by the 4p3v(M)-based and 4p3v(A)-based solvers on the correspondence is usually sufficiently accurate to score inliers. Moreover, in the LO of GC-RANSAC, this approximate model is refitted using the non-minimal 5pt solver (which is similar to the refitting that is used in ENM), and the non-minimal DLSPnP solver [dlspnp]. Similarly to ENM, the filtering (+F) that uses the correspondence does not bring an improvement that is as visible as for Poselib RANSAC. This is because for GC-RANSAC, the speedup obtained using the filtering +F is not as significant, compared to the longer running times of the LO part of GC-RANSAC.
On the other hand, the remaining two suggested modifications, i.e., the -based solvers and the refinement (+R) using the correspondence bring visible improvements. This behavior is also visible in Figure 19. Here we present an ablation study on the effects of the various modifications ( and +F/+R/+ENM, which were introduced in Sec. 3.2 of the main paper) on the 4p3v(M)-based and 4p3v(A)-based solvers. The results are reported on the St. Mary’s Church, Shop Facade, and King’s College scenes from the Cambridge Landmarks dataset [kendall2015cambridge]. These results especially highlight the importance of the refinement using the point in the third view (+R). On the other hand, the benefits of 4p3v(M)-based solvers over the 4p3v(M)-based solvers that are also visible in Table 7 are not so significant as in Poselib RANSAC. Still, 4p3v(M)-based solvers lead to improved pose accuracy.